L^{A}TEX News

Issue 12, December 1999

LPPL update

Since the release of the L^{A}TEX Project Public Licence version 1.1, we have received a small number of queries which resulted in some minor changes to improve the wording or explain the intentions better. As a consequence this release now contains LPPL 1.2 in the file lppl.txt and the previous versions as

lppl-1-0.txt and lppl-1-1.txt.

fixltx2e package

This package provides fixes to L^{A}TEX 2ε which are desirable but cannot be integrated into the L^{A}TEX 2ε kernel directly as they would produce a version incompatible to earlier releases (either in formatting or functionality).

By having these fixes in the form of a package, users can benefit from them without the danger that their documents will fail, or produce unexpected results, at other sites; this works because a document will contain a clear indication (the \usepackage line, preferably with a required date) that at least some of these fixes are required to format it.

Outcome of TUG ’99 (Vancouver)

The slides from the tug’99 presentation we gave on a new interface for L^{A}TEX class designers are available from the L^{A}TEX Project website; look for the file

tug99.pdf at:
    http://www.latex-project.org/talks/

Please note that this document was intended only to be informal “speaker’s notes” for our own use. We decided to make them available (the speaker’s notes as well as the slides that were presented) because several people requested copies after the talk. However, they are not in a polished copy-edited form and are not intended for publication.

Prototype implementations of parts of this interface are now available from:

    http://www.latex-project.org/code/experimental/

We are continuing to add new material at this location so as to stimulate further discussion of the underlying concepts. As of December 1, 1999 the following parts can be downloaded. xparse Prototype implementation of the interface for declaring document command syntax. See the .dtx files for documentation.

template Prototype implementation of the template interface (needs parts of xparse).

The file template.dtx in that directory has a large section of documentation at the front describing the commands in the interface and giving a ‘worked example’ building up some templates for caption formatting.

xcontents Interface description for table of contents data (no code yet). Coding examples have been thoroughly discussed on the latex-l list.

xfootnote Working examples for generating footnotes,

                                                   etc. Needs xparse and template.

All examples are organised in subdirectories and additionally available as gzip tar files.

Please remember that this material is intended only for experimentation and comments; thus any aspect of it, e.g., the user interface or the functionality, may change and, in fact, is very likely to change. For this reason it is explicitly forbidden to place this material on cd-rom distributions or public servers.

These concepts, as well as their implementation, are under discussion on the list LATEX-L. You can join this list, which is intended solely for discussing ideas and concepts for future versions of L^{A}TEX, by sending mail to

                                                   listserv@URZ.UNI-HEIDELBERG.DE containing the line
                                                      SUBSCRIBE LATEX-L Your Name

This list is archived and, after subscription, you can retrieve older posts to it by sending mail to the above address, containing a command such as:

                                                      GET LATEX-L LOGyymm

where yy=Year and mm=Month, e.g.

                                                      GET LATEX-L LOG9910

for all messages sent in October 1999.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1999, all rights reserved.